OpenStudy (anonymous):

(2x - 1)(x + 3) = x^2 - 3

OpenStudy (anonymous):

You should start by opening the parenthesis, and then adding like terms together: \[2x^2+6x-x-3=x^2-3 \implies 2x^2+5x-1=x^2-3 \implies x^2+5x=0\] Take x as a common factor: \[x(x+5)=0 \implies x=0, x=-5\] That's x is either 0 or -5